# Universality Diagram of Phase Transitions in Long-range Statistical Systems

### v1

@@ I Introduction
Universality lies at the heart of the renormalization-group (RG) theory of critical phenomena. Within this framework, the upper critical dimension, $d_{c}$, serves as a fundamental boundary separating distinct universality classes [PhysRevB.4.3174, PhysRevB.4.3184, PhysRevLett.28.240, chayes1987upper]. For dimensions $d<d_{c}$, critical behavior is governed by nontrivial RG fixed points, resulting in anomalous dimensions and nontrivial critical exponents. Conversely, for $d>d_{c}$, interactions become irrelevant, and the system falls into the mean-field (MF) universality class characterized by Gaussian MF exponents. At the upper critical dimension $d_{c}$, mean-field scaling laws are preserved, but modified by logarithmic corrections. Standard examples include O($n$) spin systems ($\phi^{4}$ theory) with $d_{c}=4$ [PhysRevLett.28.240] and percolation ($\phi^{3}$ theory) with $d_{c}=6$ [chayes1987upper]. Despite this well-established picture, the manifestation of universality in finite-size scaling (FSS) above $d_{c}$ remains subtle and controversial. Recent numerical studies reveal that, under periodic boundary conditions (PBCs), high-dimensional systems exhibit a characteristic two-scale FSS behavior [huang2018, fang_complete_2020, fang_logarithmic_2021, lv2021]. This phenomenon arises from the interplay between Gaussian fixed-point (GFP) scaling and complete graph (CG) scaling, where the latter corresponds to the Landau mean-field theory in a finite system. Specifically, macroscopic observables like magnetization and susceptibility are governed by CG exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})$, whereas distance-dependent quantities and correlation functions are controlled by GFP exponents $(y_{t}^{\rm GF},y_{h}^{\rm GF})$. This two-scale framework provides a crucial key to resolving long-standing debates regarding the critical scaling in high dimensions.
@@
The high-dimensional critical behavior of O(n)(n) spin models and percolation can be intuitively understood through a geometric perspective from the simple random walk (SRW) [fernandez_random_1992, zhou_random-length_2018, hara_mean-field_1990]. For O(n)(n) models, the geometric mechanism depends on the spin component $n$. In the $n\to 0$ limit, the model corresponds to the self-avoiding walk (SAW). Above $d_{c}=4$, the self-avoidance constraint becomes RG-irrelevant, and the critical two-point function recovers the SRW form $G(r)\sim|r|^{2-d}$. For O(n)(n) models with $n\geq 1$, the spin-spin correlation function maps to path weights with effective interactions determined by the intersection of trajectories [fernandez_random_1992]. Since Brownian paths have fractal dimension $2$, the intersection of two independent paths becomes marginal at $d_{c}=4$; above this, paths become mutually transparent, yielding Gaussian mean-field exponents. By analogy, high-dimensional percolation is governed by the intersection of three independent walks, which shifts the critical dimension to $d_{c}=6$ [hara_mean-field_1990]. Consequently, the SRW provides the universal geometric backbone for the mean-field physics in these systems.
@@
Complementing the thermodynamic perspective, the geometric representation of statistical systems reveals rich critical properties and offers profound insights into the original models. This is best exemplified by the Fortuin–Kasteleyn (FK) random-cluster representation [grimmett2006]. In this framework, the Ising model (equivalent to the $q=2$ Potts model) is reformulated in terms of bond configurations $A\subseteq E$ with the partition function:
@@
where $k(A)$ is the number of clusters, $v$ is the bond weight, and $|A|$ is the number of occupied bonds. Ordinary percolation is recovered as $q\to 1$, while the Ising model corresponds to $q=2$. Recent numerical and mathematical advances have uncovered that the FK–Ising model possesses two distinct upper critical dimensions [fang_geometric_2022, fang_geometric_2023, wiese_two_2024, engelenburg_one-arm_2025]. The thermodynamic upper critical dimension, $d_{c}=4$, is inherited from the $\phi^{4}$ field theory and controls standard thermodynamic observables. Meanwhile, FK clusters exhibit an additional geometric upper critical dimension, $d_{p}=6$. For $4<d<6$, the largest and second-largest clusters scale with distinct mean-field exponents derived from the CG and the GFP, respectively. In contrast, for $d\geq 6$, the FK clusters, except for the largest one, become percolation-like, characterized by strongly winding clusters and a diverging number of spanning clusters, consistent with the high-dimensional percolation model [fang_geometric_2022, fang_geometric_2023]. This double upper critical structure underscores the necessity of geometric observables for a comprehensive characterization of criticality.
@@
While spatial dimension dictates the critical behavior in short-range (SR) models, the introduction of long-range (LR) interactions provides an alternative mechanism to tune universality, adding a new layer of complexity to the landscape of critical phenomena. Since Dyson’s seminal work on the one-dimensional Ising model with algebraically decaying couplings [dysonExistencePhasetransitionOnedimensional1969], LR systems have been studied intensively [fisher1972, sak1973, luijtenBoundaryLongRangeShortRange2002, picco2012, angelini2014, defenu2023]. For a $d$-dimensional O($n$) spin model with algebraically decaying coupling strength, $J(r)\sim r^{-d-\sigma}$, the interplay between SR and LR terms generates rich critical phenomena in the $(d,\sigma)$ plane. Dimensional analysis of the effective Hamiltonian posits a threshold $\sigma_{*}=2$ separating SR and LR universality, below which the LR interaction dominates, and the system enters the nonclassical regime where critical exponents are $\sigma$-dependent. For the O($n$) spin model, this nonclassical regime extends down to $\sigma=d/2$, below which the transition is governed by the long-range Gaussian fixed point (LR-GFP), and the system is in the MF regime. The $\epsilon$ expansion analysis by Fisher suggests that the anomalous dimension follows the LR-GFP prediction, $\eta=2-\sigma$ up to $O(\epsilon^{3})$ throughout the nonclassical regime [fisher1972]. Subsequently, Sak predicted that the SR universality remains stable down to $\sigma_{*}=2-\eta_{\rm SR}$, where $\eta_{\rm SR}$ is the anomalous dimension of the corresponding SR model [sak1973]. However, recent large-scale simulations of LR-O($n$) spin models, and LR-percolation models in two dimensions have challenged this picture, providing strong evidence that the SR-LR crossover is at $\sigma_{*}=2$ [grassberger2013, LRIsing, liu_two-dimensional_2025, xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, LRHeisenberg]. These studies suggest that while the anomalous dimension $\eta$ follows the mean-field prediction $\eta=2-\sigma$ up to an extended range in the nonclassical regime, it exhibits nontrivial deviations as $\sigma\to 2^{-}$. At the same time, other critical exponents exhibit nontrivial behaviors distinct from the LR-GFP prediction, and the crossover at $\sigma_{*}=2$ may even be discontinuous, marked by a change of transition type or critical exponents [xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, LRHeisenberg, liu_two-dimensional_2025].
@@
Analogous to the SR case, the SRW offers complementary geometric insights into the LR-SR crossover. Specifically, one considers the Lévy flight – a generalization of the SRW characterized by long-range jumping probabilities decaying as $p(r)\sim 1/r^{d+\sigma}$. Regarding transport properties, the LR-SRW transitions from diffusive behavior for $\sigma>2$ to superdiffusive behavior for $\sigma\leq 2$. This signals a qualitative change of the connectivity of the underlying lattice at $\sigma=2$, which is in parallel to the boundary between SR and LR universality of LR-O($n$) models and the LR-percolation model. Furthermore, the LR-SRW describes the Goldstone mode fluctuations in the long-range ordered (LRO) phase of the LR-O(n)(n) model for $n\geq 2$. When the extensivity of a finite system is adequately taken into account, the LR-SRW correctly reproduces the scaling behaviors of Goldstone-mode fluctuations observed in LR-XY and LR-Heisenberg models for $0<\sigma\leq 2$ and $d=2$, as well as for any $\sigma>0$ and $d>2$ [LRHeisenberg]. This geometric correspondence not only reinforces the significance of the threshold at $\sigma_{*}=2$ but also implies a nontrivial connection between LR-SRW mechanics and the critical phenomenon of LR models.
@@
In parallel with these numerical breakthroughs, LR-percolation has seen significant mathematical progress [hutchcroft_dimension_2025, hutchcroft2025b, hutchcroft2025, hutchcroft2025a]. For bond percolation with connection probabilities decaying algebraically as $\sim r^{-d-\sigma}$, the asymptotics for the critical two-point function and cluster geometry have been established in various regions of the $(d,\sigma)$ plane. Notably, it has been proven that under appropriate long-range conditions, the critical two-point function behaves as $G(r)\asymp r^{-d+\sigma}$, implying $\eta=2-\sigma$ and establishing LR mean-field behavior over a wide regime. These rigorous results serve as robust anchors for constructing the LR-percolation universality diagram and provide benchmarks for conjectures in LR-O($n$) spin and FK models, for which rigorous derivations remain lacking.
@@
Motivated by these converging numerical, geometric, and mathematical insights, we revisit LR critical phenomena from a unified perspective. Our goal is to construct and compare universality diagrams in the $(d,\sigma)$ plane for LR-percolation, LR-O($n$) spin models, and the LR-FK–Ising model. We place particular emphasis on: (i) the structure of low-dimensional LR regimes; (ii) the onset of mean-field behavior and the hierarchy of upper critical dimensions; and (iii) the geometric properties of critical clusters, including the potential for additional geometric upper critical dimensions in the LR-FK–Ising case. Our construction synthesizes known SR and LR limits, the two-scale FSS above the upper critical dimension, insights from LR-SRW, high-precision numerical results in low dimensions [LRIsing, liu_two-dimensional_2025, yao2025nonclassicalregimetwodimensionallongrange], and recent rigorous mathematical results [hutchcroft_dimension_2025, hutchcroft2025b, hutchcroft2025, hutchcroft2025a]. We delineate distinct regimes of critical behavior, interpreting the roles of special boundaries to propose a coherent picture connecting percolation, O($n$), and FK–Ising models.
@@ II Finite-size Scaling Theory
Above the upper critical dimension $d_{c}$, the GFP governs the critical behavior, and the critical exponents reduce to their mean-field values. However, the FSS in this case appears to be subtle. In particular, numerical studies of high-dimensional systems with periodic boundary conditions (PBCs) reveal a characteristic two-scaling behavior, originating from the coexistence of GFP and CG asymptotics [huang2018, fang_complete_2020, lv2021, fang_logarithmic_2021, fang_geometric_2022]. The complete graph $K_{N}$ consists of $N=L^{d}$ fully connected vertices, representing the application of Landau mean-field theory to finite systems. Above $d_{c}$, the CG asymptotics account for the FSS due to PBCs and influence the scaling of macroscopic observables even though the underlying field theory is Gaussian.
@@
where $(y_{t}^{\rm GF},y_{h}^{\rm GF})$ characterize the Gaussian fixed point and $(y_{t}^{\rm CG},y_{h}^{\rm CG})$ describe the FSS behavior of the complete graph [huang2018, fang_complete_2020, lv2021, fang_logarithmic_2021, fang_geometric_2022]. A similar structure appears in the two-point correlation function, which algebraically decays to a system-size-dependent plateau,
@@
where $\hat{y}_{t}$ and $\hat{y}_{h}$ are the logarithmic correction exponents [fang_logarithmic_2021, li_logarithmic_2024]. The correlation function shows an analogous structure,
@@
Unlike O($n$) models, where logarithmic corrections arise solely from the CG sector, for percolation, the marginality of the cubic interaction induces logarithmic factors in both the GFP and CG contributions [aharony1984, kenna2006a, kenna2006, aharony1980, essam1978, ruiz-lorenzo1998].
@@
If the gyration radius is measured in an unwrapped fashion, one recovers the thermodynamic relation $C_{1}\sim R_{u}^{4}$, where $R_{u}$ is the unwrapped gyration radius [fang_geometric_2022]. The scaling of the cluster-number density, which measures the distribution of cluster size in the system, is also modified by the largest cluster size as
@@
These geometric scaling behaviors further give rise to qualitatively different topological characteristics of critical clusters in the regimes $d<6$ and $d\geq 6$. The winding number of the largest cluster can be characterized by the ratio $R_{u}/L$. For $d>6$, based on $C_{1}\sim R_{u}^{4}\sim L^{2d/3}$, the unwrapped gyration radius of the largest clusters diverges as $R_{u}\sim L^{d/6}$. Thus, for $d>6$, the largest cluster winds around the system for an increasing number of times as $L$ grows, whereas for $d<6$ the winding number remains of order unity. Furthermore, the number of spanning clusters $N_{s}$, whose unwrapped gyration radius is larger than the linear system size, shows different scalings in the two cases. For $d>6$, the number of spanning clusters diverges as $N_{s}\propto L^{d}\int_{L^{4}}n(s,L)ds\sim L^{d-6}$ [fang_geometric_2022]. However, for $d<6$, using $n(s,L)~\sim s^{-\tau}n(s/L^{d_{\mathrm{F}}})$ with $\tau=1+d/d_{\mathrm{F}}$, one finds $N_{s}\sim O(1)$ in the thermodynamic limit.

### v2

@@ I Introduction
Universality lies at the heart of the renormalization-group (RG) theory of critical phenomena. Within this framework, the upper critical dimension, $d_{c}$, serves as a fundamental boundary separating distinct universality classes [52, 53, 51, 8]. For dimensions $d<d_{c}$, critical behavior is governed by nontrivial RG fixed points, resulting in anomalous dimensions and nontrivial critical exponents. Conversely, for $d>d_{c}$, interactions become irrelevant, and the system falls into the mean-field (MF) universality class characterized by Gaussian MF exponents. At the upper critical dimension $d_{c}$, mean-field scaling laws are preserved, but modified by logarithmic corrections. Standard examples include O($n$) spin systems ($\phi^{4}$ theory) with $d_{c}=4$ [51] and percolation ($\phi^{3}$ theory) with $d_{c}=6$ [8]. Despite this well-established picture, the manifestation of universality in finite-size scaling (FSS) above $d_{c}$ remains subtle and controversial. Recent numerical studies reveal that, under periodic boundary conditions (PBCs), high-dimensional systems exhibit a characteristic two-scale FSS behavior [26, 15, 14, 42, 36, 54, 39]. This phenomenon arises from the interplay between Gaussian fixed-point (GFP) scaling and complete graph (CG) scaling, where the latter corresponds to the Landau mean-field theory in a finite system. Specifically, macroscopic observables like magnetization and susceptibility are governed by CG exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})$, whereas distance-dependent quantities and correlation functions are controlled by GFP exponents $(y_{t}^{\rm GF},y_{h}^{\rm GF})$. This two-scale framework provides a crucial key to resolving long-standing debates regarding the critical scaling in high dimensions.
@@
The high-dimensional critical behavior of O(n)(n) spin models and percolation can be intuitively understood through a geometric perspective from the simple random walk (SRW) [18, 60, 24]. For O(n)(n) models, the geometric mechanism depends on the spin component $n$. In the $n\to 0$ limit, the model corresponds to the self-avoiding walk (SAW). Above $d_{c}=4$, the self-avoidance constraint becomes RG-irrelevant, and the critical two-point function recovers the SRW form $G(r)\sim|r|^{2-d}$. For O(n)(n) models with $n\geq 1$, the spin-spin correlation function maps to path weights with effective interactions determined by the intersection of trajectories [18]. Since Brownian paths have fractal dimension $2$, the intersection of two independent paths becomes marginal at $d_{c}=4$; above this, paths become mutually transparent, yielding Gaussian mean-field exponents. By analogy, high-dimensional percolation is governed by the intersection of three independent walks, which shifts the critical dimension to $d_{c}=6$ [24]. Consequently, the SRW provides the universal geometric backbone for the mean-field physics in these systems.
@@
Complementing the thermodynamic perspective, the geometric representation of statistical systems reveals rich critical properties and offers profound insights into the original models. This is best exemplified by the Fortuin–Kasteleyn (FK) random-cluster representation [23]. In this framework, the Ising model (equivalent to the $q=2$ Potts model) is reformulated in terms of bond configurations $A\subseteq E$ with the partition function:
@@
where $k(A)$ is the number of clusters, $v$ is the bond weight, and $|A|$ is the number of occupied bonds. Ordinary percolation is recovered as $q\to 1$, while the Ising model corresponds to $q=2$. Recent numerical and mathematical advances have uncovered that the FK–Ising model possesses two distinct upper critical dimensions [16, 17, 50, 12]. The thermodynamic upper critical dimension, $d_{c}=4$, is inherited from the $\phi^{4}$ field theory and controls standard thermodynamic observables. Meanwhile, FK clusters exhibit an additional geometric upper critical dimension, $d_{p}=6$. For $4<d<6$, the largest and second-largest clusters scale with distinct mean-field exponents derived from the CG and the GFP, respectively. In contrast, for $d\geq 6$, the FK clusters, except for the largest one, become percolation-like, characterized by strongly winding clusters and a diverging number of spanning clusters, consistent with the high-dimensional percolation model [16, 17]. This double upper critical structure underscores the necessity of geometric observables for a comprehensive characterization of criticality.
@@
While spatial dimension dictates the critical behavior in short-range (SR) models, the introduction of long-range (LR) interactions provides an alternative mechanism to tune universality, adding a new layer of complexity to the landscape of critical phenomena. Since Dyson’s seminal work on the one-dimensional Ising model with algebraically decaying couplings [11], LR systems have been studied intensively [19, 46, 41, 44, 4, 10]. For a $d$-dimensional O($n$) spin model with algebraically decaying coupling strength, $J(r)\sim r^{-d-\sigma}$, the interplay between SR and LR terms generates rich critical phenomena in the $(d,\sigma)$ plane. Dimensional analysis of the effective Hamiltonian posits a threshold $\sigma_{*}=2$ separating SR and LR universality, below which the LR interaction dominates, and the system enters the nonclassical regime where critical exponents are $\sigma$-dependent. For the O($n$) spin model, this nonclassical regime extends down to $\sigma=d/2$, below which the transition is governed by the long-range Gaussian fixed point (LR-GFP), and the system is in the MF regime. The $\epsilon$ expansion analysis by Fisher suggests that the anomalous dimension follows the LR-GFP prediction, $\eta=2-\sigma$ up to $O(\epsilon^{3})$ throughout the nonclassical regime [19]. Subsequently, Sak predicted that the SR universality remains stable down to $\sigma_{*}=2-\eta_{\rm SR}$, where $\eta_{\rm SR}$ is the anomalous dimension of the corresponding SR model [46]. However, recent large-scale simulations of LR-O($n$) spin models, and LR-percolation models in two dimensions have challenged this picture, providing strong evidence that the SR-LR crossover is at $\sigma_{*}=2$ [22, 55, 38, 56, 58, 57]. These studies suggest that while the anomalous dimension $\eta$ follows the mean-field prediction $\eta=2-\sigma$ up to an extended range in the nonclassical regime, it exhibits nontrivial deviations as $\sigma\to 2^{-}$. At the same time, other critical exponents exhibit nontrivial behaviors distinct from the LR-GFP prediction, and the crossover at $\sigma_{*}=2$ may even be discontinuous, marked by a change of transition type or critical exponents [56, 58, 57, 38].
@@
Analogous to the SR case, the SRW offers complementary geometric insights into the LR-SR crossover. Specifically, one considers the Lévy flight – a generalization of the SRW characterized by long-range jumping probabilities decaying as $p(r)\sim 1/r^{d+\sigma}$. Regarding transport properties, the LR-SRW transitions from diffusive behavior for $\sigma>2$ to superdiffusive behavior for $0<\sigma\leq 2$. This signals a qualitative change of the connectivity of the underlying lattice at $\sigma=2$, which is in parallel to the boundary between SR and LR universality of LR-O($n$) models and the LR-percolation model. Furthermore, the LR-SRW describes the Goldstone mode fluctuations in the long-range ordered (LRO) phase of the LR-O(n)(n) model for $n\geq 2$. When the extensivity of a finite system is adequately taken into account, the LR-SRW correctly reproduces the scaling behaviors of Goldstone-mode fluctuations observed in LR-XY and LR-Heisenberg models for $0<\sigma\leq 2$ and $d=2$, as well as for any $\sigma>0$ and $d>2$ [57]. This geometric correspondence not only reinforces the significance of the threshold at $\sigma_{*}=2$ but also implies a nontrivial connection between LR-SRW mechanics and the critical phenomenon of LR models.
@@
In parallel with these numerical breakthroughs, LR-percolation has seen significant mathematical progress [30, 27, 28, 29]. For bond percolation with connection probabilities decaying algebraically as $\sim r^{-d-\sigma}$, the asymptotics for the critical two-point function and cluster geometry have been established in various regions of the $(d,\sigma)$ plane. Notably, it has been proven that under appropriate long-range conditions, the critical two-point function behaves as $G(r)\asymp r^{-d+\sigma}$, implying $\eta=2-\sigma$ and establishing LR mean-field behavior over a wide regime. These rigorous results serve as robust anchors for constructing the LR-percolation universality diagram and provide benchmarks for conjectures in LR-O($n$) spin and FK models, for which rigorous derivations remain lacking.
@@
Motivated by these converging numerical, geometric, and mathematical insights, we revisit LR critical phenomena from a unified perspective. Our goal is to construct and compare universality diagrams in the $(d,\sigma)$ plane for LR-percolation, LR-O($n$) spin models, and the LR-FK–Ising model. We place particular emphasis on: (i) the structure of low-dimensional LR regimes; (ii) the onset of mean-field behavior and the hierarchy of upper critical dimensions; and (iii) the geometric properties of critical clusters, including the potential for additional geometric upper critical dimensions in the LR-FK–Ising case. Our construction synthesizes known SR and LR limits, the two-scale FSS above the upper critical dimension, insights from LR-SRW, high-precision numerical results in low dimensions [55, 38, 58], and recent rigorous mathematical results [30, 27, 28, 29]. We delineate distinct regimes of critical behavior, interpreting the roles of special boundaries to propose a coherent picture connecting percolation, O($n$), and FK–Ising models.
@@ II Finite-size Scaling Theory
Above the upper critical dimension $d_{c}$, the GFP governs the critical behavior, and the critical exponents reduce to their mean-field values. However, the FSS in this case appears to be subtle. In particular, numerical studies of high-dimensional systems with periodic boundary conditions (PBCs) reveal a characteristic two-scaling behavior, originating from the coexistence of GFP and CG asymptotics [26, 15, 42, 14, 16]. The complete graph $K_{N}$ consists of $N=L^{d}$ fully connected vertices, representing the application of Landau mean-field theory to finite systems. Above $d_{c}$, the CG asymptotics account for the FSS due to PBCs and influence the scaling of macroscopic observables even though the underlying field theory is Gaussian.
@@
where $(y_{t}^{\rm GF},y_{h}^{\rm GF})$ characterize the Gaussian fixed point and $(y_{t}^{\rm CG},y_{h}^{\rm CG})$ describe the FSS behavior of the complete graph [26, 15, 42, 14, 16]. A similar structure appears in the two-point correlation function, which algebraically decays to a system-size-dependent plateau,
@@
where $\hat{y}_{t}$ and $\hat{y}_{h}$ are the logarithmic correction exponents [14, 35]. The correlation function shows an analogous structure,
@@
Unlike O($n$) models, where logarithmic corrections arise solely from the CG sector, for percolation, the marginality of the cubic interaction induces logarithmic factors in both the GFP and CG contributions [2, 33, 32, 3, 13, 45].
@@
If the gyration radius is measured in an unwrapped fashion, one recovers the thermodynamic relation $C_{1}\sim R_{u}^{4}$, where $R_{u}$ is the unwrapped gyration radius [16]. The scaling of the cluster-number density, which measures the distribution of cluster size in the system, is also modified by the largest cluster size as
@@
These geometric scaling behaviors further give rise to qualitatively different topological characteristics of critical clusters in the regimes $d<6$ and $d\geq 6$. The winding number of the largest cluster can be characterized by the ratio $R_{u}/L$. For $d>6$, based on $C_{1}\sim R_{u}^{4}\sim L^{2d/3}$, the unwrapped gyration radius of the largest clusters diverges as $R_{u}\sim L^{d/6}$. Thus, for $d>6$, the largest cluster winds around the system for an increasing number of times as $L$ grows, whereas for $d<6$ the winding number remains of order unity. Furthermore, the number of spanning clusters $N_{s}$, whose unwrapped gyration radius is larger than the linear system size, shows different scalings in the two cases. For $d>6$, the number of spanning clusters diverges as $N_{s}\propto L^{d}\int_{L^{4}}n(s,L)ds\sim L^{d-6}$ [16]. However, for $d<6$, using $n(s,L)~\sim s^{-\tau}n(s/L^{d_{\mathrm{F}}})$ with $\tau=1+d/d_{\mathrm{F}}$, one finds $N_{s}\sim O(1)$ in the thermodynamic limit.